Let  A = { n ∈ Z ∣ n ≡ 1 ( mod 2 ), then  A   is the set of

Group of answer choices

even integers

odd integers

Z∖{0}

Z

None of the above

Q2

If A = { 1 } and  B = { 2 }, then the power set,  P ( A × B ) is

Group of answer choices

{ϕ,{A×B}}

{ϕ,{1×2}}

{ϕ,{(1,2)}}

{ϕ,{A}×{B}}

None of the above
Answer #1

A = \{ n \epsilon Z :\;n \equiv 1 ( mod \;2 ),
\\ This \;expression\;implies\;\\we\;get\;remainder\;1\;if\;divided\;by\;2\\
Therefore n is 2k + 1 form number
\boldsymbol{{\color{Red} That\;is \;n\;an\;odd\;number}}

A\times B=cartesian\;product= \{\,(a,b)\mid a\in A\ {\mbox{ and }}\ b\in B\,\}
Ax B= (1,2)
Power of the (AxB) is {0, 1, 2, (1, 2)}
\\\left \{ A \right \}=\left \{ \phi ,1 \right \}\;\;and\left \{ B \right \}=\left \{ \phi, 2 \right \}\\ \left \{ \left \{ A \right \}\times\left \{ B \right \} \right \}=\left \{ \phi ,1,2,(1,2) \right \}
{\color{Red} \boldsymbol{\{\phi ,\{A\}x\{B\}\}}}

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image